# Medium Access for Push-Pull Data Transmission in 6G Wireless Systems

### v1

@@ Abstract
Index Terms:
IoT connectivity, pull-based communication, push-based communication, medium access control, timing, data relevance
@@ I Introduction
The novel services envisioned with 6G systems expect flexible communication protocols designed to accommodate different applications and requirements. The network slicing approaches proposed for 5G can create multiple virtual networks tailored to specific service needs on a shared physical infrastructure to meet the requirements of different traffic classes. However, due to the dynamic variability of end-to-end goals over time as well as the use of AI by the communicating parties, 6G systems will require agile and robust Medium Access Control (MAC) protocols to accommodate diverse application-driven requirements, not catered efficiently by the existing network slicing strategies [1]. The communication protocols for 6G systems should be designed considering the growing intelligence in the communicating nodes and the coexistence of diverse application requirements and communication modalities.
The resulting high reliability and flexibility requirements pose two key challenges: first, where to process the information under strict latency-reliability requirements, and secondly, when and how to collect the data for processing. The former aspect has been extensively discussed in the domain of edge computing and task offloading. The latter aligns with the goal-oriented communication paradigm [3], which optimizes communication decisions based on the Value of Information (VoI) of updates: when transmitting all available data is impossible due to timing or energy constraints, goal-oriented schemes privilege information that is both novel and relevant for the receiver, i.e., is both not directly predictable from previously available information and useful for the task at hand [2].
@@
Figure 1: A toy example of an IoT scenario where a BS collects measurements from IoT devices to construct a digital twin (DT) of the system. The coexistence of push- and pull-based communication enables the BS to coordinate requests for the most informative data to update the DT model while allowing IoT devices to immediately report detected anomalies.
@@
The design of protocols and MAC mechanisms for the coexistence of push- and pull-based traffic supporting goal-oriented Key Performance Indicators (KPIs), such as estimation error or control performance, in addition to traditional timing metrics, is a key enabler for real-time applications such as DTs and XR [8]; addressing this gap is imperative to the design of efficient 6G systems. As knowledge about the environment is itself distributed, centralizing decision-making at the BS or distributing it to the sensors leads to highly different information structures and signaling requirements.
This article introduces the design criteria for MAC protocols to accommodate push- and pull-based communications, offering guidelines for organizing intelligent operation in 6G wireless access networks. To that end, we design a time-division mechanism and provide numerical studies in selected use cases to demonstrate the validity of the developed analytical models and their potential for application-driven 6G network operation.
@@ III-A Centralized decision-making
CFC-pull/push is a frame structure similar to the contention-free and contention-based mechanism in WiFi technology. Each frame’s organization is notified to the devices through control signaling [10]: firstly, the BS decides the device scheduling for the pull-based communication phase, taking into account the available pull-slots, and notifies devices by broadcasting the scheduling information. This unicast scheduling enables the BS to retrieve sensor readings from specific devices at specific times without packet collisions. Conversely, during the push-contention phase, each device transmits an update – if it deems it is important enough – by contending for the channel with other push-based traffic. The major challenge is to design the resource allocation scheme that maximizes the goal-oriented utility for both phases [10]. Fig. 4 shows the average incoming traffic that can be served by the system with a target success probability as a function of $\alpha$. Every point in the plot represents a different number of slots reserved for responding to pull requests, $\lfloor\alpha S\rfloor$. As expected, increasing this value leads to a decrease in the acceptable push traffic, while accommodating a higher pull traffic, and vice versa. Naturally, reducing the target success probability allows the BS to manage a higher overall load.
[…]
Figure 4: CFC-pull/push achievable incoming traffic satisfying a target transmission success probability $p_{\mathrm{th}}$ as a function of $\alpha$ [10].
@@ III-B Decentralized intelligence
In the decentralized intelligence scenario, devices make local observations and estimate their expected value to decide when to transmit. As individual devices have limited knowledge about others’ observations, transmissions are subject to collisions, and thus to the loss of timely and accurate status updates. Devices then need to decide whether and when to attempt a transmission so as to satisfy the KPIs for the goal-oriented tasks, such as VoI [5]. As even the BS has partial information on the value of data at the nodes, this leads to a different choice of MAC frame design and access protocols, as the frame structure needs to be adapted significantly to match the requirements set by the emerging applications and use cases.
@@
For instance, consider a distributed learning setting for collective intelligence. The BS would use private data available at the IoT devices, which have limited computational and communication capabilities, to train high-quality learning models. To do so, the BS can reuse the CFC-pull/push frame structure and the corresponding communication protocol to solicit application-relevant data. This means devising efficient scheduling policies to accelerate high-quality training, where contention-free pull-reserved slots are allocated to the devices with the most useful information in each frame. A wrong attribution of the value of data from specific devices during scheduling, due to the partial knowledge of the BS, can result in a weaker model with poor generalization performance. To overcome this issue, devices that are not accommodated in the pull slots need transmission opportunities to share a better representation of their data. This can be done by arranging available communication resources for pull- and push-based slots, such that a high-quality learning model can be trained in a communication-efficient manner, as shown in Fig. 6. A reference protocol design for such a case is defined in [12].
@@
Furthermore, devices need to balance the urgency of updates that might have to wait for a long time before the next pull-based slot with the relative scarcity of push-based slots and the consequent risk of collisions. Goal-oriented operation might require reasoning about other devices’ beliefs, exploiting publicly available information to come to a consensus without further signaling. Naturally, aside from the computational load that this puts on the nodes, it also requires a careful design of the beacon and acknowledgment messages.
5102030400.880.880.880.880.90.90.90.90.920.920.920.920.940.940.940.94Number of push IoT devices per frameTest AccuracyRandomStrategic push-pullOnly-pull
@@
Figure 6: Learning performance under a fixed number of slots per frame; strategic access push-pull improves accuracy over Only-pull, and outperforms Random scheduling significantly.
@@ IV-A Intelligent MAC protocols
However, the risk of missing out unforeseen critical updates, such as alarms, always exists in the absence of available push communication resources. In practice, the previous considerations require the BS to be able to dynamically modify the MAC structure to achieve different trade-offs, switching between CFC-pull/push and RCSC-pull/push frames and adjusting the number of pull, push, or shared slots (see Fig. 3). This also translates to the need for specific and reliable control signals broadcasting the information regarding the MAC design to all devices in a timely manner. In the extreme case, the MAC structure could be modified on a frame-by-frame basis, but trade-offs between frame flexibility and the overhead caused by the additional control information should be investigated.
@@ IV-B Energy-Efficient Connectivity
While employing WuR can improve the overall energy efficiency, it also adds design challenges. Highly informative feedback from the BS, for example, aiming to improve future transmission policy through strategic push, cannot be realistically performed through WuSs due to their low data rate [14]. Furthermore, a dynamic adoption of WuSs within the flexible MAC protocols is needed to tailor the correct push-pull paradigm per the design space, accordingly to the current context of data collection and the goal-oriented scenario (see Fig. 2). Such dynamic adaptation should consider the pros and cons of grant-based and grant-free access in terms of system-level performance, including energy efficiency.
@@
A seamless integration of push-pull data transmissions in future 6G networks necessitates both a higher-layer application able to translate the current goal-oriented requirements and the status of the network into a viable allocation of network resources, and advanced Radio Resource Control (RRC) mechanisms capable of dynamic, real-time radio resource allocation based on the chosen MAC structure. The former can be obtained through the development of Open-Radio Access Network (O-RAN)-compliant xApps/rApps [15], while the latter requires flexible management of the physical (PHY) channels.
@@
xApps operating within the Near-Real-Time RAN Intelligent Controller (Near-RT RIC) can serve as proxies between the entity requesting the data, e.g., the DT, and the next-generation Node B (gNB). On one side, xApps can interpret high-level requirements arriving through the Y1 interface – such as data content preferences and deadlines – and combine this information with network status feedback provided by the gNB via the E2 interface [15]. This enables the conversion of goal-oriented requirements into resource configurations, which are sent back to the gNB for implementation through dynamic RRC. Moreover, xApps are responsible for processing data acquired from the network before transmitting it to the end destination via the Y1 interface. Finally, in scenarios where multiple agents request data with potentially conflicting objectives, an xApp must aggregate and prioritize these requirements to ensure fair and efficient management of network resources. Similar operations can be implemented in rApps operating within the Non-Real-Time RIC (Non-RT RIS) [15], where near-real-time control loops are not needed.
@@
To effectively implement the resource configurations determined by xApps, the gNB must employ RRC to dynamically manage radio resource allocation for each PHY channel, organizing their placement within the MAC frame. In 5G, contention-based communication utilizes the Physical Random Access Channel (PRACH), primarily for initial access. This channel can support contention-based transmission in push-pull MAC designs, i.e., for the push portion of the CFC-pull/push frame and throughout the RCSC-pull/push frame. Conversely, the Physical Uplink Shared Channel (PUSCH) handles scheduled uplink data transmissions, enabling uni-cast pull-based communications within the CFC-pull/push. Moreover, the Physical Downlink Control Channel (PDCCH) can convey control signaling regarding dynamic RRC information, including MAC design decisions and scheduling directives. A flexible allocation of these channels across available PHY resources is crucial to accommodate diverse push-pull data transmission paradigms. Moreover, integrating energy-efficient technologies such as WuR introduces additional complexity in control signaling interactions. For example, to enable the CFC-pull/push frame structure with UCWu, a WuS containing the device ID could be sent via the PDCCH– or a newly defined physical WuS channel – before each pull slot, which in turn utilizes the PUSCH. Then, the PRACH would be employed to implement push slots, potentially occurring with a periodicity lower than $10$ ms – the minimum allowed by the 5G standard – depending on the frame duration [10]. Therefore, exploring more flexible RRC approaches is essential for integrating push-pull coexistence into future 6G networks.
@@ V Conclusion
Shashi Raj Pandey is currently an assistant professor at the Connectivity Section, Aalborg University, Denmark. He received his Ph.D. in computer science and engineering from Kyung Hee University, South Korea, in 2021. His research interests include network economics, wireless communications and networking, distributed machine learning, and semantic communications.
Fabio Saggesse is a postdoctoral researcher in the Connectivity Section at Aalborg University, Denmark. He earned his Ph.D. in Information Engineering from the University of Pisa, Italy, in 2022. In 2024, he received the Marie Skłodowska-Curie Postdoctoral Fellowship. His research focuses on PHY/MAC wireless connectivity, goal-oriented communications, and integrated sensing and communications.
@@
Federico Chiariotti is currently an assistant professor at the University of Padova, Italy, where he also received his PhD in information engineering in 2019. He has authored over 100 peer-reviewed papers on latency optimization, semantic and goal-oriented communication, and machine learning in networking, and is currently an Associate Editor of IEEE TWC and IEEE Net. Lett.

### v2

@@ I Introduction
The novel services envisioned with 6G systems expect flexible communication protocols designed to accommodate different applications and requirements. The network slicing approaches proposed for 5G can create multiple virtual networks tailored to specific service needs on a shared physical infrastructure to meet the requirements of different traffic classes. However, due to the dynamic variability of end-to-end goals over time as well as the use of AI by the communicating parties, 6G systems will require agile and robust Medium Access Control (MAC) protocols to accommodate diverse application-driven requirements, not catered efficiently by the existing network slicing strategies [1]. As such, the MAC protocols should, by design, adhere to goal-oriented communication approaches, which is missing in the related works on protocols. The communication protocols for 6G systems should be designed considering the growing intelligence in the communicating nodes and the coexistence of diverse application requirements and communication modalities.
The resulting high reliability and flexibility requirements pose two key challenges: first, where to process the information under strict latency-reliability requirements, and secondly, when and how to collect the data for processing. The former aspect has been extensively discussed in the domain of edge computing and task offloading. The latter aligns with the goal-oriented communication paradigm [3], which optimizes communication decisions based on the Value of Information (VoI) of updates: when transmitting all available data is impossible due to timing or energy constraints, goal-oriented schemes privilege information that is both novel and relevant for the receiver, i.e., is both not directly predictable from previously available information and useful for the task at hand [4, 5].
@@
Figure 1: A toy example of an IoT scenario where a BS collects measurements from IoT devices to construct a DT of the system. The coexistence of push- and pull-based communication enables the BS to coordinate requests for the most informative data to update the DT model while allowing IoT devices to immediately report detected anomalies.
@@
The design of protocols and MAC mechanisms for the coexistence of push- and pull-based traffic supporting goal-oriented Key Performance Indicators (KPIs), such as estimation error or control performance, in addition to traditional timing metrics, is a key enabler for real-time applications such as DTs and XR [9]; addressing this gap is imperative to the design of efficient 6G systems. As knowledge about the environment is itself distributed, centralizing decision-making at the BS or distributing it to the sensors leads to highly different information structures and signaling requirements. Our previous work [10, 11] has begun exploring some of these challenges, but this paper provides a systematic treatment of integrated access framework, considering the fact that AI-driven operation expands the design space of medium access protocols.
This article introduces the design criteria for future MAC protocols to accommodate push- and pull-based communications, offering guidelines for organizing intelligent operation in 6G wireless access networks, as explained in sections II and III. To this end, we provide a time-division frame structure that can accommodate these novel schemes, as well as analyzing their potential for application-driven 6G network operation in section IV.
@@ III-A Centralized decision-making
CFC-pull/push is a frame structure similar to the contention-free and contention-based mechanism in WiFi technology. Each frame’s organization is notified to the devices through control signaling [10]: firstly, the BS decides the device scheduling for the pull-based communication phase, taking into account the available pull-slots, and notifies devices by broadcasting the scheduling information. This unicast scheduling enables the BS to retrieve sensor readings from specific devices at specific times without packet collisions. Conversely, during the push-contention phase, each device transmits an update – if it deems it is important enough – by contending for the channel with other push-based traffic. The major challenge is to design the resource allocation scheme that maximizes the goal-oriented utility for both phases [10]. Fig. 4 shows the maximum average incoming traffic that can be handled by the system when constraining the successful delivery of packets within a target latency $L$ with reliability $99\%$. Simulations are made according to the following setting: a 10 ms frame comprises a total of $S=100$ slots, with pull and push packets occupying 5 and 1 slots each, respectively, due to the different data carried [10]; a 1-persistent retransmission scheme in case of push collision is employed; similarly, unscheduled pull packets are rescheduled in the next frame. Every point in the plot represents a different percentage of the slots reserved for pull requests, $\alpha$. Any value above the curves cannot satisfy the latency constraint, giving a measure of scalability of the system with respect to the incoming system load. As expected, increasing this value leads to a decrease in the acceptable push traffic, while accommodating a higher pull traffic, and vice versa. Naturally, reducing the target latency allows the BS to manage a higher overall load.
00.511.522.53012345510152025303540455055606570758085909551015202530354045505560657075808590955101520253035404550556065707580859095Avg. achievable incoming pull traffic [k-packets/s] Avg. achievable incoming push traffic [k-packets/s] $\alpha$ ($\%$)$L=20$ msL=30L=30 msL=50L=50 ms
Figure 4: CFC-pull/push achievable incoming traffic satisfying a target latency $L$ with reliability 99% as a function of $\alpha$ [10].
@@
The proposed frame designs are general and allow the use of any scheduler and random access approach tailored for managing transmissions within contention-free and contention-based slots, respectively. More advanced contention resolution techniques – e.g., by making the use of successive interference cancellation – can be integrated to enhance the access performance in contention-based sub-frames, though the quantification of potential gains requires further investigation.
@@ III-B Decentralized intelligence
In the decentralized intelligence scenario, devices make local observations and estimate their expected value to decide when to transmit. As individual devices have limited knowledge about others’ observations, transmissions are subject to collisions, and thus to the loss of timely and accurate status updates. Devices then need to decide whether and when to attempt a transmission so as to satisfy the KPIs for the goal-oriented tasks, such as VoI [4]. As even the BS has partial information on the value of data at the nodes, the access protocols must adapt to accommodate push- and pull-based traffic depending on the requirements set by the emerging applications and use cases, exploiting the flexibility in the proposed frame structure. Additionally, devices might also have limited computational and storage capabilities: in these settings, the decision-making needs to be shifted towards the BS, but centrally computed policies that can be executed locally may still provide good performance.
@@
For instance, consider a distributed learning setting for collective intelligence. The BS would use private data available at the IoT devices, which have limited computational and communication capabilities, to train high-quality learning models. To do so, the BS can reuse the CFC-pull/push frame structure and the corresponding communication protocol to solicit application-relevant data. This means devising efficient scheduling policies to accelerate high-quality training, where contention-free pull-reserved slots are allocated to the devices with the most useful information in each frame. A wrong attribution of the value of data from specific devices during scheduling, due to the partial knowledge of the BS, can result in a weaker model with poor generalization performance. To overcome this issue, devices that are not accommodated in the pull slots need (push) transmission opportunities to autonomously share a better representation of their data. This can be done by arranging available communication resources for both pull- and push-based slots, such that a high-quality learning model can be trained in a communication-efficient manner, as shown in Fig. 6. The figure demonstrates how proposed strategic push-pull approach outperforms Random, Centralized, and Only-pull approaches in terms of test accuracy, while providing competitive accuracy to the Oracle scheme, where the full information on data availability and timing is exploited. A reference protocol design for such a case is defined in [13].
@@
Furthermore, devices need to balance the urgency of updates that might have to wait for a long time before the next pull-based slot with the relative scarcity of push-based slots and the consequent risk of collisions. Goal-oriented operation might require inferences about other devices’ beliefs, exploiting publicly available information to come to a consensus without further signaling. Naturally, aside from the computational load that this puts on the nodes, it also requires a careful design of the beacon and acknowledgment messages. However, a full analysis of the additional energy consumption this imposes on the devices is strongly dependent on the mechanism or AI tool that they use to estimate value and make decisions, and is beyond the scope of this paper.
5102030400.880.880.90.90.920.920.940.94Number of push UEs NNTest accuracyOracleRandomStrategic push-pullCentralizedOnly-pull
@@
Figure 6: Learning performance under a fixed number of slots per frame; strategic access push-pull improves accuracy over Only-pull and Centralized approach, where contention-free collection of fraction of data is done. The method stays competitive with the Oracle and outperforms Random scheduling significantly.
@@ IV-A Intelligent MAC protocols
However, the risk of missing out unforeseen critical updates, such as alarms, always exists in the absence of available push communication resources. In practice, the previous considerations require the BS to be able to dynamically modify the MAC structure to achieve different trade-offs, switching between CFC-pull/push and RCSC-pull/push frames and adjusting the number of pull, push, or shared slots (see Fig. 3). This also translates to the need for specific and reliable control signals broadcasting the information regarding the MAC design to all devices in a timely manner. In the extreme case, the MAC structure could be modified on a frame-by-frame basis, but trade-offs between frame flexibility and the overhead caused by the additional control information should be investigated when considering the system-level performance of grant-based and grant-free access.
@@ IV-B Energy-Efficient Connectivity
While employing WuR can improve the overall energy efficiency, it also adds design challenges. Highly informative feedback from the BS, e.g., aiming to improve future transmission policy through strategic push, cannot be realistically performed through WuSs due to their low data rate [14]. Furthermore, a dynamic adoption of WuSs within the flexible MAC protocols is needed to tailor the correct push-pull paradigm per the design space, accordingly to the current context of data collection and the goal-oriented scenario (see Fig. 2). An energy efficient implementation of such dynamic MAC reconfiguration requires a thorough investigation of the energy costs associated with signaling overhead.
@@
A seamless integration of a push-pull approach in future 6G networks necessitates both a higher-layer application able to translate the current goal-oriented requirements and the status of the network into a viable allocation of network resources, and advanced Radio Resource Control (RRC) mechanisms capable of dynamic, real-time radio resource allocation based on the chosen MAC structure. The former can be obtained through the development of Open-Radio Access Network (O-RAN)-compliant xApps/rApps [15], while the latter requires flexible management of the physical (PHY) channels.
@@
xApps operating within the Near-Real-Time RAN Intelligent Controller can serve as proxies between the entity requesting the data, e.g., the DT, and the next-generation Node B (gNB). xApps can interpret high-level requirements arriving through the Y1 interface – such as data content preferences and deadlines – and combine this information with network status feedback provided by the gNB via the E2 interface [15]. This enables the conversion of goal-oriented requirements into resource configurations, to be implemented through dynamic RRC. Moreover, xApps can process data acquired from the network before the transmission to the end destination via the Y1 interface. Finally, in scenarios where multiple network agents potentially have conflicting objectives, an xApp must aggregate and prioritize these requirements to ensure efficient resource allocations. This allocation problem is highly dependent on the scenario and the needs of the involved applications, requiring a complex multi-objective optimization which is in itself an open challenge. When near-real-time is not required, similar operations can be implemented in rApps within the Non-Real-Time RIC [15].
@@
To effectively implement the resource configurations determined by xApps, the RRC must dynamically manage radio resource allocation for each PHY channel, organizing their placement within the MAC frame. In 5G, access is performed over the Physical Random Access Channel (PRACH). This channel can support contention-based transmission in push-pull MAC designs, i.e., for the push portion of the CFC-pull/push frame and throughout the RCSC-pull/push frame. Conversely, the Physical Uplink Shared Channel (PUSCH) handles scheduled transmissions, enabling uni-cast pull-based communications within the CFC-pull/push. Moreover, the Physical Downlink Control Channel (PDCCH) can convey control signaling regarding dynamic RRC information, including MAC frame design and scheduling directives. A flexible allocation of these channels across available PHY resources is crucial to accommodate diverse push-pull paradigms. Moreover, integrating energy-efficient technologies such as WuR introduces additional complexity in control signaling interactions. For example, to enable the CFC-pull/push frame with UCWus, a WuS containing the device ID could be sent via the PDCCH– or a newly defined physical WuS channel – before each pull slot employing the PUSCH. Then, the PRACH would implement push slots, potentially occurring with a periodicity lower than $10$ ms – the minimum allowed by the 5G standard – depending on the frame duration [10]. Thus, exploring more flexible RRC approaches is essential for integrating push-pull coexistence into future 6G networks.
@@ V Conclusion
Shashi Raj Pandey is currently an assistant professor at the Connectivity Section, Aalborg University, Denmark. He received his Ph.D. in computer science and engineering from Kyung Hee University, South Korea, in 2021. His research interests include network economics, wireless communications and networking, distributed machine learning, and semantic communications, and is currently an Associate Editor of IEEE TMLCN.
Fabio Saggese is currently a Marie Skłodowska-Curie Postdoctoral Fellow at the Department of Information Engineering, University of Pisa, where he received his Ph.D. in 2022. His research focuses on PHY/MAC wireless connectivity, goal-oriented communications, and integrated sensing and communications.
@@
Federico Chiariotti is an assistant professor at the University of Padova, Italy, where he also received his PhD in information engineering in 2019. He has authored over 100 peer-reviewed papers on latency optimization and semantic and goal-oriented communication and is currently an Associate Editor of IEEE TWC and IEEE Net. Lett.